OpenStudy (anonymous):

what is the probability of rolling a 5 on a six sided die and then rolling a 2? 1.1/36 2.1/4 3.1/5 4.5/6

OpenStudy (anonymous):

(1/6)*(1/6)

OpenStudy (anonymous):

P(5)=1/6 P(2)=1/6 P(5,2)=(1/6)(1/6)=(1/36)

OpenStudy (anonymous):

Thanks